Henry invests $2432 in a savings account for 3 years that earns 1.4% compounded quarterly.


How much interest will he have earned at the end of 3 years?


1
Expert's answer
2021-05-07T09:58:48-0400

A=P{1+r/100}^12

r=1.4÷4=0.35

Therefore

r=0.35÷100=0.0035

A=$2432 {1+0.0035}^12

A=$2432 {1.0035}^12

A=$2432 {1.0428}

A=$2536.13

I=A-P

I=$2536.13-$2432

I=$104.13